# MADAR: Efficient Continual Learning for Malware Analysis with Diversity-Aware Replay

### v1

@@ VI-A MADAR
Building on the exploratory analysis in Section V, we postulate that stratified sampling, where replay samples are chosen based on their representation in malware families, may better preserve the model’s stability versus random sampling as used in GRS. Moreover, we also seek to capture the diversity within each family’s data distribution. We do this by selecting a balance between representative samples that are close to each other and anomalous samples that are separated. While any single anomalous sample is not as important to learn and remember as a single representative sample, a collection of anomalous samples helps to track the diversity within a class.
@@
2init   $\mathcal{P}$; init   $\mathcal{D}\leftarrow\{M_{f}:M_{c}\}$;
3if $c=0$ then
@@ 10       $X_{good},X_{mal}\leftarrow\mathcal{P}$;
12      $\beta_{A}$, $\beta_{S}\leftarrow\beta_{M}~{}.~{}\Omega$;
13      if $\Psi$ then
@@ 17             $\mathcal{F_{MC}}\leftarrow X_{f}$;
19                   $\mathcal{MC}\leftarrow|\mathcal{D}|$;$\mathcal{B_{F}}\leftarrow(\mathcal{F_{MC}}/\mathcal{MC})*\beta_{M}$;
20
@@ 33      $Y_{train}\leftarrow(\texttt{concat}((Y_{c},Y_{replay}))$;
For each family $f$, we set its family budget $\mathcal{B}_{f}$— the number of samples to select from $f$—using two sub-sampling variants: Ratio budgeting and Uniform budgeting.
@@ VI-B MADARθ
[…]
Adaptation of Batch Normalization.  During the forward pass, batch normalization is omitted in MADARθ to preserve the diversity of the activation distributions. While batch normalization typically stabilizes learning and improves generalization, it can homogenize activations, reducing the diversity essential for identifying anomalies. This adaptation enhances the performance of MADARθ, as confirmed by empirical results.
@@ VII Evaluation
Baselines Joint 96.4$\pm$0.3
None 93.1$\pm$0.1
GRS 93.6$\pm$0.3 94.1$\pm$1.3 95.3$\pm$0.2 95.3$\pm$0.7 95.9$\pm$0.1 95.8$\pm$0.6 96.0$\pm$0.3
Prior Work ER [24] 80.6$\pm$0.1 73.5$\pm$0.5 70.5$\pm$0.3 69.9$\pm$0.1 70.0$\pm$0.1 70.0$\pm$0.1 70.0$\pm$0.1
AGEM [25] 80.5$\pm$0.1 73.6$\pm$0.2 70.4$\pm$0.3 70.0$\pm$0.1 70.0$\pm$0.2 70.0$\pm$0.1 70.0$\pm$0.1
GR [17] 93.1$\pm$0.2
RtF [41] 93.2$\pm$0.2
BI-R [14] 93.4$\pm$0.1
MADAR MADAR-R 93.7$\pm$0.1 94.7$\pm$0.1 95.4$\pm$0.1 95.3$\pm$0.6 96.0$\pm$0.1 96.1$\pm$0.1 96.1$\pm$0.1
MADAR-U 93.6$\pm$0.2 94.0$\pm$0.2 95.1$\pm$0.1 95.3$\pm$0.1 95.5$\pm$0.1 95.7$\pm$0.1 95.8$\pm$0.1
MADARθ-R 93.6$\pm$0.1 94.4$\pm$0.3 95.3$\pm$0.2 95.8$\pm$0.1 96.1$\pm$0.1 96.1$\pm$0.1 96.1$\pm$0.1
MADARθ-U 93.5$\pm$0.2 94.1$\pm$0.2 94.9$\pm$0.1 95.2$\pm$0.2 95.6$\pm$0.1 95.7$\pm$0.1 95.7$\pm$0.1
(a) MADAR Ratio
(b) MADAR Uniform
(c) MADARθ Ratio
(d) MADARθ Uniform
Figure 7: EMBER Domain-IL: Comparison of the MADAR-R, MADAR-U, MADARθ-R, and MADARθ-U with Joint baseline.
@@
Baselines Joint 97.3$\pm$0.1
None 94.4$\pm$0.1
GRS 95.3$\pm$0.1 96.4$\pm$0.1 96.9$\pm$0.1 97.1$\pm$0.1 97.1$\pm$0.1 97.2$\pm$0.1 97.2$\pm$0.1
Prior Work ER [24] 40.4$\pm$0.1 40.1$\pm$0.1 41.1$\pm$0.2 42.6$\pm$0.1 44.0$\pm$0.1 45.9$\pm$0.1 48.6$\pm$1.1
AGEM [25] 45.4$\pm$0.1 47.4$\pm$0.2 49.2$\pm$0.2 53.7$\pm$0.6 54.2$\pm$0.3 54.8$\pm$0.4 56.7$\pm$0.3
GR [17] 93.3$\pm$0.4
RtF [41] 93.4$\pm$0.2
BI-R [14] 93.5$\pm$0.1
MADAR MADAR-R 95.8$\pm$0.1 96.6$\pm$0.1 96.9$\pm$0.1 97.0$\pm$0.1 97.0$\pm$0.1 97.0$\pm$0.1 97.0$\pm$0.1
MADAR-U 95.7$\pm$0.1 95.5$\pm$0.1 95.2$\pm$0.2 95.2$\pm$0.1 95.4$\pm$0.1 95.8$\pm$0.2 96.3$\pm$0.2
MADARθ-R 95.8$\pm$0.2 96.6$\pm$0.1 96.9$\pm$0.1 96.9$\pm$0.1 97.1$\pm$0.1 97.1$\pm$0.1 97.2$\pm$0.1
MADARθ-U 95.6$\pm$0.1 96.1$\pm$0.1 96.6$\pm$0.1 96.8$\pm$0.1 97.0$\pm$0.1 97.1$\pm$0.1 97.1$\pm$0.1
(a) MADAR Ratio
(b) MADAR Uniform
(c) MADARθ Ratio
(d) MADARθ Uniform
Figure 8: AZ Domain-IL: Comparison of the MADAR-R, MADAR-U, MADARθ-R, and MADARθ-U with Joint baseline.
@@ VII-A Domain-IL
In EMBER, we have 12 tasks, each representing the monthly data distribution spanning January–December 2018. Our results, detailed in Table II, provide a comprehensive view of each method’s performance, reported as the average accuracy over all tasks $\mathbf{\overline{AP}}$. Additionally, Figure 7 illustrates the progression of average accuracy over time compared to the Joint baseline.
The informal lower and upper performance bounds for this configuration are approximated by the None and Joint methods, achieving $\mathbf{\overline{AP}}$ scores of 93.1% and 96.4%, respectively. Meanwhile, GRS serves as a strong baseline, providing unbiased sampling without incorporating sample diversity awareness.
At a lower budget of 1K, MADAR-R, MADAR-U, and MADARθ-R exhibit competitive performance, all achieving $\mathbf{\overline{AP}}$ of over $93.6$%, significantly outperforming prior approaches. This highlights their ability to effectively utilize limited resources. In particular, MADAR-R achieves the highest accuracy at this budget, with $\mathbf{\overline{AP}}$ of $93.7\%$.
As the memory budget increases, the performance of all MADAR and MADARθ variants improves steadily. At a budget of 200K, MADAR-R and MADARθ-R achieve an impressive $\mathbf{\overline{AP}}$ of $96.0\%$ and $96.1\%$, respectively, closely approaching the $96.4\%$ achieved by the Joint baseline, which utilizes over 670K samples. Uniform strategies, including MADAR-U and MADARθ-U, are only slightly behind, with $\mathbf{\overline{AP}}$ values of $95.5\%$ and $95.6\%$, respectively.
@@
Baselines Joint 86.5$\pm$0.4
None 26.5$\pm$0.2
GRS 51.9$\pm$0.4 70.3$\pm$0.5 75.4$\pm$0.7 82.0$\pm$0.2 83.5$\pm$0.1 84.3$\pm$0.3 84.6$\pm$0.2
Prior Work TAMiL [13] 32.2$\pm$0.3 33.1$\pm$0.2 35.3$\pm$0.2 36.7$\pm$0.1 38.2$\pm$0.3 37.2$\pm$0.2 38.8$\pm$0.2
iCaRL [15] 53.9$\pm$0.7 58.7$\pm$0.7 60.0$\pm$1.0 63.9$\pm$1.2 64.6$\pm$0.8 65.5$\pm$1.0 66.8$\pm$1.1
ER [24] 27.5$\pm$0.1 27.8$\pm$0.1 28.0$\pm$0.1 27.9$\pm$0.1 28.0$\pm$0.1 28.0$\pm$0.1 28.2$\pm$0.1
AGEM [25] 27.3$\pm$0.1 27.4$\pm$0.1 27.7$\pm$0.1 28.5$\pm$0.1 28.2$\pm$0.1 28.3$\pm$0.1 28.2$\pm$0.1
GR [17] 26.8$\pm$0.2
RtF [41] 26.5$\pm$0.1
BI-R [14] 26.9$\pm$0.1
MalCL [11] 54.5$\pm$0.3
MADAR MADAR-R 68.0$\pm$0.4 73.6$\pm$0.2 76.0$\pm$0.3 81.5$\pm$0.2 83.2$\pm$0.2 83.8$\pm$0.2 84.0$\pm$0.2
MADAR-U 66.4$\pm$0.4 76.5$\pm$0.2 79.4$\pm$0.4 83.8$\pm$0.2 84.8$\pm$0.1 85.5$\pm$0.1 85.8$\pm$0.3
MADARθ-R 67.9$\pm$0.3 72.7$\pm$0.5 72.7$\pm$0.5 81.7$\pm$0.2 83.2$\pm$0.1 83.9$\pm$0.1 84.5$\pm$0.2
MADARθ-U 67.5$\pm$0.3 76.4$\pm$0.4 78.5$\pm$0.4 84.1$\pm$0.1 85.3$\pm$0.1 85.8$\pm$0.2 86.2$\pm$0.2
@@
For the experiments with AZ-Domain, we consider 9 tasks, each representing a yearly data distribution from 2008 to 2016. The performance of each method is presented in Table III as $\mathbf{\overline{AP}}$ and compared to two baselines: None, which achieves $94.4\%$, and Joint, which reaches $97.3\%$. Additionally, Figure 8 illustrates the progression of average accuracy across tasks, highlighting the comparison with the Joint baseline.
@@
At higher budgets, ranging from 100K to 400K, MADAR-R continues to demonstrate high $\mathbf{\overline{AP}}$ scores of up to $97.0\%$, closely matching GRS and only marginally below the Joint baseline, which requires significantly more training samples (680K). Notably, MADARθ-R exhibits comparable performance, reaching a peak $\mathbf{\overline{AP}}$ of $97.2\%$ at the highest budget level, further underscoring the efficacy of our diversity-aware approach.
In summary, these results empirically demonstrate the effectiveness of MADAR’s diversity-aware sample selection in enhancing the efficiency and accuracy of malware classification in Domain-IL scenarios. MADAR-R and MADARθ-R, in particular, consistently either yield on-par or outperform GRS while delivering significant improvements over prior methods.
@@
Baselines Joint 94.2$\pm$0.1
None 26.4$\pm$0.2
GRS 43.8$\pm$0.7 62.9$\pm$0.8 70.2$\pm$0.4 83.0$\pm$0.3 86.4$\pm$0.2 88.2$\pm$0.2 89.1$\pm$0.2
Prior Work TAMiL [13] 53.4$\pm$0.3 55.2$\pm$0.3 57.6$\pm$0.3 60.8$\pm$0.2 63.5$\pm$0.1 65.3$\pm$0.5 67.7$\pm$0.3
iCaRL [15] 43.6$\pm$1.2 54.9$\pm$1.0 61.7$\pm$0.7 77.2$\pm$0.4 81.5$\pm$0.6 83.4$\pm$0.5 84.6$\pm$0.5
ER [24] 50.8$\pm$0.7 58.3$\pm$0.6 58.9$\pm$0.2 59.2$\pm$0.8 62.9$\pm$0.7 63.1$\pm$0.5 64.2$\pm$0.4
AGEM [25] 27.3$\pm$0.7 28.0$\pm$1.4 27.1$\pm$0.3 28.0$\pm$0.6 28.2$\pm$1.0 29.8$\pm$2.6 28.0$\pm$0.8
GR [17] 22.7$\pm$0.3
RtF [41] 22.9$\pm$0.3
BI-R [14] 23.4$\pm$0.2
MalCL [11] 59.8$\pm$0.4
MADAR MADAR-R 59.4$\pm$0.6 67.8$\pm$0.9 71.9$\pm$0.5 82.9$\pm$0.2 86.3$\pm$0.1 88.2$\pm$0.2 89.1$\pm$0.1
MADAR-U 57.3$\pm$0.5 70.4$\pm$0.4 76.2$\pm$0.2 86.8$\pm$0.1 89.8$\pm$0.1 91.0$\pm$0.1 91.5$\pm$0.1
MADARθ-R 58.8$\pm$0.3 66.2$\pm$0.7 71.0$\pm$0.7 81.2$\pm$0.3 85.1$\pm$0.2 86.9$\pm$0.2 88.1$\pm$0.1
MADARθ-U 58.5$\pm$0.7 70.1$\pm$0.2 74.7$\pm$0.2 85.5$\pm$0.1 88.7$\pm$0.1 90.3$\pm$0.2 90.7$\pm$0.1
@@ VII-B Class-IL
At higher budgets, MADAR-U and MADARθ-U continue to excel, with MADARθ-U achieving the best results overall. At a 20K budget, MADARθ-U reaches an $\mathbf{\overline{AP}}$ of $86.2\%$, nearly equaling the Joint baseline, which uses over 150 times more training samples. These results clearly demonstrate the effectiveness of MADAR’s diversity-aware sample selection and the effectiveness of MADAR-U and MADARθ-U in leveraging limited resources.
@@
In summary, our experiments demonstrate the effectiveness of MADAR’s diversity-aware replay techniques in the Class-IL setting for both EMBER and AZ datasets. While GRS shows significant improvement with larger budgets, MADAR’s uniform variants consistently outperform it across all budget levels. These results underscore MADAR’s ability to mitigate catastrophic forgetting and enhance malware classification performance, even in resource-constrained environments.
@@
Baselines Joint 97.0$\pm$0.3
None 74.6$\pm$0.7
GRS 86.9$\pm$0.3 87.4$\pm$0.3 93.6$\pm$0.3 94.4$\pm$0.2 94.7$\pm$0.3 94.9$\pm$0.1 95.0$\pm$0.1
Prior Work TAMiL [13] 72.8$\pm$0.1 81.5$\pm$0.3 86.9$\pm$0.2 88.1$\pm$0.3 90.3$\pm$0.1 93.2$\pm$0.3 94.2$\pm$0.7
ER [24] 67.4$\pm$0.3 84.9$\pm$0.2 89.5$\pm$0.5 93.9$\pm$0.2 94.8$\pm$0.2 95.2$\pm$0.1 95.4$\pm$0.1
AGEM [25] 79.6$\pm$0.2 81.7$\pm$0.2 83.8$\pm$0.4 84.9$\pm$0.2 86.1$\pm$0.2 88.9$\pm$0.2 89.3$\pm$0.1
GR [17] 79.8$\pm$0.3
RtF [41] 77.8$\pm$0.2
BI-R [14] 87.2$\pm$0.3
MADAR MADAR-R 92.1$\pm$0.2 92.3$\pm$0.9 93.8$\pm$0.2 94.2$\pm$0.1 94.8$\pm$0.2 95.7$\pm$0.2 95.6$\pm$0.1
MADAR-U 93.4$\pm$0.2 93.7$\pm$0.3 93.9$\pm$0.3 94.8$\pm$0.2 95.6$\pm$0.1 95.7$\pm$0.1 95.8$\pm$0.2
MADARθ-R 93.1$\pm$0.2 93.3$\pm$0.1 93.6$\pm$0.1 94.3$\pm$0.1 94.6$\pm$0.2 94.8$\pm$0.2 94.7$\pm$0.3
MADARθ-U 93.2$\pm$0.1 93.1$\pm$0.2 93.8$\pm$0.2 94.4$\pm$0.1 94.8$\pm$0.1 95.3$\pm$0.2 95.5$\pm$0.3
@@ VII-C Task-IL
Baselines Joint 98.8$\pm$0.2
None 74.5$\pm$0.2
GRS 85.2$\pm$0.1 89.2$\pm$0.2 90.8$\pm$0.1 91.6$\pm$0.2 93.5$\pm$0.1 93.9$\pm$0.1 95.2$\pm$0.1
Prior Work TAMiL 80.5$\pm$0.4 85.3$\pm$0.6 91.5$\pm$0.2 92.1$\pm$0.1 93.5$\pm$0.1 94.0$\pm$0.2 94.8$\pm$0.2
ER 83.6$\pm$0.2 90.2$\pm$0.1 92.3$\pm$0.3 95.6$\pm$0.1 96.2$\pm$0.1 96.8$\pm$0.2 97.5$\pm$0.2
AGEM 76.7$\pm$0.5 82.8$\pm$0.2 85.3$\pm$0.1 85.6$\pm$0.2 86.7$\pm$0.2 88.9$\pm$0.2 91.3$\pm$0.3
GR 75.6$\pm$0.2
RtF 74.2$\pm$0.3
BI-R 85.4$\pm$0.2
MADAR MADAR-R 86.0$\pm$0.3 90.3$\pm$0.2 92.4$\pm$0.1 95.8$\pm$0.2 96.7$\pm$0.1 97.1$\pm$0.1 97.9$\pm$0.2
MADAR-U 88.1$\pm$0.3 92.9$\pm$0.2 94.5$\pm$0.3 97.2$\pm$0.2 98.1$\pm$0.1 98.5$\pm$0.1 98.7$\pm$0.1
MADARθ-R 87.3$\pm$0.3 90.6$\pm$0.2 93.2$\pm$0.2 95.7$\pm$0.2 95.9$\pm$0.1 96.6$\pm$0.1 96.9$\pm$0.1
MADARθ-U 87.9$\pm$0.2 90.8$\pm$0.2 93.6$\pm$0.1 96.2$\pm$0.3 97.2$\pm$0.2 97.5$\pm$0.2 98.1$\pm$0.1
@@ VII-D Analysis and Discussion
Our results demonstrate that MADAR yields markedly better performances compared to previous methods for both the EMBER and AZ datasets across all CL settings. This clearly indicates that diversity-aware replay is effective in preserving the stability of a CL-based system for malware classification, while prior CL techniques largely fail to achieve acceptable performance.
@@
GRS remains a strong baseline at high budgets.  While MADAR consistently outperforms GRS in low-resource settings, GRS performs comparably at higher budgets, particularly in Domain-IL. This suggests that diversity-aware replay is most impactful when the number of available samples per class is limited, whereas uniform selection provides sufficient representation at larger budgets.

### v2

@@ VI-A MADAR
Building on the exploratory analysis in Section V, we postulate that stratified sampling, where replay samples are chosen based on their representation in malware families, may better preserve the model’s stability versus random sampling as used in GRS. Moreover, we also seek to capture the heterogeneity within each family’s data distribution. We do this by selecting a balance between representative samples that are close to each other and anomalous samples that are separated. While any single anomalous sample is not as important to learn and remember as a single representative sample, a collection of anomalous samples helps to track the heterogeneity within a class.
@@
21ex
3init   $\mathcal{P}$; init   $\mathcal{D}\leftarrow\{M_{f}:M_{c}\}$;
4if $c=0$ then
@@ 11  $X_{good},X_{mal}\leftarrow\mathcal{P}$;
13 $\beta_{A}$, $\beta_{S}\leftarrow\beta_{M}~.~\Omega$; if $\Psi$ then
@@ 18     $\mathcal{F_{MC}}\leftarrow X_{f}$;
20        $\mathcal{MC}\leftarrow|\mathcal{D}|$$\mathcal{B_{F}}\leftarrow(\mathcal{F_{MC}}/\mathcal{MC})*\beta_{M}$;
21
@@ 34 $Y_{train}\leftarrow(\texttt{concat}((Y_{c},Y_{replay}))$;
For each family $f$, we set its family budget $\mathcal{B}_{f}$— the number of samples to select from ff—using two sub-sampling variants: Ratio budgeting and Uniform budgeting.
@@ VI-B MADARθ
Selection of the Layer $\mathcal{L}$.  The choice of $\mathcal{L}$ is critical in MADARθ to ensure that feature representations are captured without interference from the model’s final classification stage. Empirical testing revealed that the penultimate layers, denoted as act4 for the EMBER dataset and act5 for the AZ datasets, provide optimal results such that $\mathcal{L}_{\text{EMBER}}=\texttt{act4},\quad\mathcal{L}_{\text{AZ}}=\texttt{act5}.$
Adaptation of Batch Normalization.  During the forward pass, batch normalization is omitted in MADARθ to preserve the heterogeneity of the activation distributions. While batch normalization typically stabilizes learning and improves generalization, it can homogenize activations, reducing the heterogeneity essential for identifying anomalies. This adaptation enhances the performance of MADARθ, as confirmed by empirical results.
@@ VII Evaluation
[…]
@@
[…]
@@ VII-A Domain-IL
In EMBER, we have 12 tasks, each representing the monthly data distribution spanning January–December 2018. Our results, detailed in Table II, provide a comprehensive view of each method’s performance, reported as the average accuracy over all tasks $\mathbf{\overline{AP}}$.
The informal lower and upper performance bounds for this configuration are approximated by the None and Joint methods, achieving $\mathbf{\overline{AP}}$ scores of 93.1% and 96.4%, respectively. Meanwhile, GRS serves as a strong baseline, providing unbiased sampling without incorporating sample heterogeneity awareness.
At a lower budget of 1K, MADAR-R, MADAR-U, and MADARθ-R exhibit competitive performance, all achieving $\mathbf{\overline{AP}}$ of over 93.693.6%, significantly outperforming prior approaches. This highlights their ability to effectively utilize limited resources. In particular, MADAR-R achieves the highest accuracy at this budget, with $\mathbf{\overline{AP}}$ of $93.7\%$. As the memory budget increases, the performance of all MADAR and MADARθ variants improves steadily. At a budget of 200K, MADAR-R and MADARθ-R achieve an impressive $\mathbf{\overline{AP}}$ of $96.0\%$ and $96.1\%$, respectively, closely approaching the $96.4\%$ achieved by the Joint baseline, which utilizes over 670K samples. Uniform strategies, including MADAR-U and MADARθ-U, are only slightly behind, with $\mathbf{\overline{AP}}$ values of $95.5\%$ and $95.6\%$, respectively.
@@
[…]
@@
For the experiments with AZ-Domain, we consider 9 tasks, each representing a yearly data distribution from 2008 to 2016. The performance of each method is presented in Table III as $\mathbf{\overline{AP}}$ and compared to two baselines: None, which achieves $94.4\%$, and Joint, which reaches $97.3\%$.
@@
At higher budgets, ranging from 100K to 400K, MADAR-R continues to demonstrate high $\mathbf{\overline{AP}}$ scores of up to $97.0\%$, closely matching GRS and only marginally below the Joint baseline, which requires significantly more training samples (680K). Notably, MADARθ-R exhibits comparable performance, reaching a peak $\mathbf{\overline{AP}}$ of $97.2\%$ at the highest budget level, further underscoring the efficacy of our distribution-aware approach.
In summary, these results empirically demonstrate the effectiveness of MADAR’s distribution-aware sample selection in enhancing the efficiency and accuracy of malware classification in Domain-IL scenarios. MADAR-R and MADARθ-R, in particular, consistently either yield on-par or outperform GRS while delivering significant improvements over prior methods.
@@
[…]
@@ VII-B Class-IL
At higher budgets, MADAR-U and MADARθ-U continue to excel, with MADARθ-U achieving the best results overall. At a 20K budget, MADARθ-U reaches an $\mathbf{\overline{AP}}$ of $86.2\%$, nearly equaling the Joint baseline, which uses over 150 times more training samples. These results clearly demonstrate the effectiveness of MADAR’s distribution-aware sample selection and the effectiveness of MADAR-U and MADARθ-U in leveraging limited resources.
@@
In summary, our experiments demonstrate the effectiveness of MADAR’s distribution-aware replay techniques in the Class-IL setting for both EMBER and AZ datasets. While GRS shows significant improvement with larger budgets, MADAR’s uniform variants consistently outperform it across all budget levels. These results underscore MADAR’s ability to mitigate catastrophic forgetting and enhance malware classification performance, even in resource-constrained environments.
@@
[…]
@@ VII-C Task-IL
[…]
@@ VII-D Analysis and Discussion
Our results demonstrate that MADAR yields markedly better performances compared to previous methods for both the EMBER and AZ datasets across all CL settings. This clearly indicates that distribution-aware replay is effective in preserving the stability of a CL-based system for malware classification, while prior CL techniques largely fail to achieve acceptable performance.
@@
GRS remains a strong baseline at high budgets.  While MADAR consistently outperforms GRS in low-resource settings, GRS performs comparably at higher budgets, particularly in Domain-IL. This suggests that distribution-aware replay is most impactful when the number of available samples per class is limited, whereas uniform selection provides sufficient representation at larger budgets.